OpenStudy (anonymous):

Find the zeros of the polynomial function and state the multiplicity of each. f(x) = 5(x + 8)2(x - 8)3

OpenStudy (campbell_st):

well all you need for the zeros is to solve x - 8 = 0 and x + 8 = 0 multiplicity is the number of times the factor occurs... so look at the powers of each factor

OpenStudy (jdoe0001):

\(\bf f(x) = 5(x + 8)^2(x - 8)^3\\ \textit{setting y = 0, we get}\\ 0 = 5(x + 8)^2(x - 8)^3\\ \textit{solve for }\quad x \quad \textit{keep in mind that}\\ 5(x + 8)^2(x - 8)^3\implies 5(x + 8)(x + 8)(x - 8)(x - 8)(x - 8)\)